OpenStudy (andriod09):

Guys, if you're factoring, and you have a problem that has \[(r^0)5\] there wouldn't be a r, would there??

OpenStudy (anonymous):

Maybe it will help if you post the whole problem, but mathematically r^0=1, so that is a number.

OpenStudy (andriod09):

The original problem is this: \[(-2r^0s^5)^5\]

OpenStudy (jdoe0001):

anyhow, ANY VALUE raised to 0 is 1, ANY

OpenStudy (andriod09):

So then it would be \[(-2r^1s^5)^5\]

OpenStudy (jdoe0001):

anyhow, ANY VALUE raised to 0 is \(\huge \color{red}1\), ANY

OpenStudy (andriod09):

... I understand. Is that what i asked though? No. I asked if that was correct, which I have yet to be answered. >.<

OpenStudy (jdoe0001):

hehe

OpenStudy (jdoe0001):

I never said it was raised to 1, I said IT WAS 1

OpenStudy (andriod09):

I'm not solving for vaiables though. i'm factoring. >.< "factoring"

OpenStudy (jdoe0001):

same applies, say r=2, 2^0=1, say r=1,000,000,000, 000, 1,000,000,000, 000^2=1

OpenStudy (andriod09):

Makes no sense.

OpenStudy (jdoe0001):

wooops, I meant 1,000,000,000, 000^0 btw :)

OpenStudy (anonymous):

no, there would be any r and any value (except 0) raised to 0 =1

OpenStudy (andriod09):

so is my problem where the r is correct?

OpenStudy (jdoe0001):

dunno, what do you think is \(\large r^{0*5}\)?

OpenStudy (anonymous):

(r^0)^5 = 1^5=1